# Understanding the temperature response of biological systems: Part II - Network-level mechanisms and emergent dynamics

### v1

@@ Introduction
In Part I of this review [1], we surveyed phenomenological models that can describe biological temperature response curves, together with microscopic single-reaction level theories that derive temperature dependence from physical and chemical principles. While these approaches clarify how temperature affects individual processes and observed rate curves, they do not capture how system-level temperature responses emerge from interactions among multiple reactions and regulatory interactions.
Here, in part II of this series, we focus on network-level mechanisms that generate emergent temperature dependence in biological systems. In such models, temperature not only modulates individual reaction rates, but also the collective dynamics of interconnected pathways, feedback loops, and multistep processes. We consider two complementary classes of network-level descriptions. Deterministic models use coupled differential equations to describe how temperature shapes the dynamics of biochemical networks, while stochastic models (for instance formulated as Markov chains) capture the role of fluctuations, multi-step transitions, and statistical structure in determining system-level timing and rates.
Together, these network-level frameworks show how simple Arrhenius-like temperature dependence at the level of individual reactions can give rise to complex, non-Arrhenius temperature response curves at the system level. By explicitly linking empirical temperature responses to network organization and dynamical structure, these models provide mechanistic insight into robustness, temperature compensation, and the constraints governing biological timing across scales.
@@
A minimal and well-characterized example of a temperature-sensitive biochemical oscillator is the early embryonic cell cycle of the frog Xenopus laevis. Entry into mitosis (M phase) and exit back into interphase are controlled by the kinase Cdk1, which becomes active only when bound to Cyclin B and appropriately (de)phosphorylated. Active Cyclin B–Cdk1 drives mitotic events and promotes Cyclin degradation, completing a negative feedback loop.
The core regulatory architecture therefore consists of Cyclin B synthesis, its Cdk1-dependent degradation, and a bistable switch controlling Cdk1 activation (Fig. 1A). Together, the slow accumulation and degradation of Cyclin and the fast switching of Cdk1 form a classic relaxation oscillator. In its simplest formulation, this system can be described by a two-variable ODE model for Cyclin and active Cdk1 [2, 3]. The model reproduces the characteristic sawtooth-like Cyclin dynamics and sharp Cdk1 activation pulses seen in extracts (Fig. 1B). In the phase plane, the Cdk1 activation module forms an S-shaped nullcline, and its intersection with the Cyclin nullcline gives rise to an unstable state (US) around which a limit cycle emerges (Fig. 1C).
To investigate temperature scaling, Arrhenius laws can be assigned to the kinetic rates governing Cyclin synthesis ($k_{s}$), degradation ($k_{d}$), and Cdk1 activation and inactivation. Even this minimal two-ODE system is sufficient to explain non-Arrhenius period scaling. When individual reactions have different activation energies (Fig. 1D), the ratio $k_{s}/k_{d}$ changes with temperature, shifting the Cyclin nullcline in the phase plane. If $k_{s}$ increases more steeply with temperature than $k_{d}$, the Cyclin nullcline moves upward. At high enough temperatures the nullclines cease to intersect on the middle branch of the Cdk1 nullcline, the US disappears, and the limit cycle collapses into a high-Cdk1 fixed point (persistent M phase). Conversely, at low temperatures the system becomes trapped in a low-Cdk1 interphase-like state. These bifurcations generate thermal limits and a curved, non-Arrhenius period–temperature relation (Fig. 1D).
Experimental measurements in cycling Xenopus egg extracts support this mechanistic picture [3]. Within the viable embryonic temperature range, the rising (interphase) and falling (M-phase) segments of the oscillation scale differently with temperature, consistent with Cyclin synthesis being more temperature-sensitive than its degradation. This difference largely explains the non-Arrhenius scaling of the total period in the physiological range.
@@
Early mechanistic proposals by Hastings and Sweeney suggested that opposing biochemical reactions within the clock could have counterbalancing temperature responses [9]. This idea motivated a series of mathematical models in which different parts of the transcription–translation feedback loop scale differently with temperature [10, 11, 12]. However, perfect compensation of temperature dependencies is not robust to parameter variation or genetic perturbations.
@@
Both the embryonic cell cycle oscillator and circadian models have a similar network architecture, combining negative feedback with additional regulatory steps that change the effective feedback strength [14]. However, the consequences for temperature responses differ. In the circadian model, the interconversion between protein states occurs on timescales comparable to the temperature-sensitive synthesis and degradation reactions. This allows redistribution between isoforms to buffer temperature-induced changes and thus stabilize the period. In the embryonic cell cycle oscillator, by contrast, the dominant slow process is Cyclin accumulation, whereas switching between Cdk1 states is very fast. This strong timescale separation means that changes in the relative temperature sensitivities of synthesis and degradation directly shift the nullclines and hence the period, decreasing internal buffering. Thus, despite similarities in network structure, differences in relative timescales and in how feedback is routed lead to very different temperature-response phenotypes.
@@
Recent work has established a statistical framework for the generic temperature dependence of large biochemical networks which are stochastic and for which each rate follows an Arrhenius law [15]. We note that even at the level of individual reactions, the Arrhenius form derived from Kramers’ theory relies on the assumption of sufficiently large energy barriers. When barriers become comparable to thermal fluctuations, this approximation breaks down and non-Arrhenius scaling can arise already at the single-reaction level, as shown using mean first-passage time (MFPT) approaches [17]. Starting from the master equation of a time-continuous, space-discrete Markov process, the analysis uses the graph-theoretic representation of the MFPT to a target state [18]. The overall rate $r$ is the inverse MFPT, and its temperature dependence $\ln{r(T)}$ can be expanded in a Taylor series whose coefficients are related to the distributions of the total activation energies along the spanning trees $E_{\mathcal{T}}$ and spanning forests of two trees $E_{\mathcal{F}}$ of the stochastic network (see Fig. 2) according to:
@@
where $\langle E\rangle$ and $\sigma^{2}$ denote the means and variances of the respective energy distributions. This prediction agrees well with simulations of random networks and published developmental-rate data in flies which have previously been fitted by the phenomenological quadratic exponential described in Part I [19, 20, 21]. This limit requires that the difference in the activation energies along the trees and the forests stays remains distinct in the large network limit, typically assured by a bias in the activation energies for rates towards the target state. Moreover, it is only fully valid if the network is sufficiently complex to have a macroscopic number of spanning trees and forests.
@@
The master equation approach has also been applied to a specific, and analytically tractable, network topology: a linear chain of $n$ reversible reactions (Fig. 2 C) [16]. Such cascades provide a reasonable representation of many multi-step biochemical processes. However, since they have only one spanning tree they do not fulfill the requirements leading to the result in (2). Still, the analysis of this model reveals that, even in this case, the generic quadratic exponential arises. Additionally, it shows how and why this quadratic scaling can break down at extreme temperatures.
Each individual step follows an Arrhenius law with a randomly distributed activation energy. At the reference temperature $T_{0}$, the cascade is assumed to be forward-biased, meaning that forward rates are, on average, larger than backward rates. Computing the MFPT from the initial to the final state shows that the overall temperature dependence of the rate falls into three regimes.
In the physiologically relevant range around $T_{0}$, the many forward transitions collectively dominate the MFPT. By the law of large numbers, the sum of their activation energies approaches a normal distribution, and the rate obeys a quadratic exponential scaling (2) with parameters
$E^{*}=\langle E\rangle,\qquad B=\frac{\sigma_{E}^{2}}{2},\qquad$ $r^{*}=\frac{n}{\langle 1/r_{f}^{*}\rangle}.$ (3)
@@
At low temperatures, however, the system undergoes a sudden transition. As backward reactions become comparably faster, the system spends an increasing amount of time in cycles of forward and backward reactions before reaching the final state. Below a critical temperature $T_{+}$, a single cycle dominates the MFPT. In this regime, the rate scales Arrhenius-like with a large positive activation energy that depends on the sum and difference of the activation energies within this critical cycle.
A similar breakdown of the quadratic exponential occurs at high temperatures. Above a second critical temperature $T_{-}$, one cycle again dominates the MFPT leading to an overall rate that follows the Arrhenius law. However, in this regime the effective activation energy becomes negative. The critical temperatures separating the three regimes are:
@@
Biological processes exhibit diverse temperature responses across scales, yet many of their large-scale features arise from shared physical and organizational principles. In Part I of this review [1], we showed how phenomenological descriptions and microscopic reaction-level theories capture robust regularities in rate–temperature relationships and clarify why deviations from simple Arrhenius behavior are common. Here, in Part II, we have focused on how these local temperature dependencies are transformed by network organization into emergent system-level temperature responses.
@@ Acknowledgements
The work is supported by grants from Internal funds KU Leuven (C14/23/130, LG). JBV thanks the German Academic Scholarship Foundation (Studienstiftung des Deutschen Volkes) for support.

### v2

@@ Abstract
Figure 1: Temperature influences biological systems across scales and can be described using different classes of models. (A) Conceptual illustration of Arrhenius and non-Arrhenius behavior. In an Arrhenius plot (log rate versus $1/T$), simple reactions follow a straight line, whereas biological processes typically show curvature and thermal limits ($T_{\min},T_{\max}$). (B) Overview of modeling frameworks used to describe temperature responses, organized by level of description. Phenomenological models provide empirical fits to observed rate–temperature curves, while microscopic models derive rate–temperature relationships from reaction-level kinetics. At a higher level, network-level models—either deterministic or stochastic—capture how temperature affects coupled biochemical or regulatory systems. The phenomenological and microscopic approaches form the focus of Part I [1], whereas deterministic and stochastic network models are the focus of this work in Part II [2].
@@ Introduction
In Part I of this review [1], we surveyed phenomenological models that can describe biological temperature response curves, together with microscopic single-reaction level theories that derive temperature dependence from physical and chemical principles (Figure 1A-B). While these approaches clarify how temperature affects individual processes and observed rate curves, they do not capture how system-level temperature responses emerge from interactions among multiple reactions and regulatory interactions.
Here, in part II of this series, we focus on network-level mechanisms that generate emergent temperature dependence in biological systems (Figure 1B). In such models, temperature not only modulates individual reaction rates, but also the collective dynamics of interconnected pathways, feedback loops, and multistep processes. We consider two complementary classes of network-level descriptions. Deterministic models use coupled differential equations to describe how temperature shapes the dynamics of biochemical networks, while stochastic models (for instance formulated as Markov chains) capture the role of fluctuations, multi-step transitions, and statistical structure in determining system-level timing and rates.
Together, these network-level frameworks show how simple Arrhenius-like temperature dependence at the level of individual reactions can give rise to complex, non-Arrhenius temperature response curves at the system level (Figure 1A). By explicitly linking empirical temperature responses to network organization and dynamical structure, these models provide mechanistic insight into robustness, temperature compensation, and the constraints governing biological timing across scales.
@@
A minimal and well-characterized example of a temperature-sensitive biochemical oscillator is the early embryonic cell cycle of the frog Xenopus laevis. Entry into mitosis (M phase) and exit back into interphase are controlled by the kinase Cdk1, which becomes active only when bound to Cyclin B and appropriately (de)phosphorylated. Active Cyclin B–Cdk1 drives mitotic events and promotes Cyclin B degradation, completing a negative feedback loop.
The core regulatory architecture therefore consists of Cyclin B synthesis, its Cdk1-dependent degradation, and a bistable switch controlling Cdk1 activation (Fig. 2A). Together, the slow accumulation and degradation of Cyclin B and the fast switching of Cdk1 form a classic relaxation oscillator. In its simplest formulation, this system can be described by a two-variable ODE model for Cyclin B and active Cdk1 [3, 4]. The model reproduces the characteristic sawtooth-like Cyclin B dynamics and sharp Cdk1 activation pulses seen in extracts (Fig. 2B). In the phase plane, the Cdk1 activation module forms an S-shaped nullcline, and its intersection with the Cyclin B nullcline gives rise to an unstable state (US) around which a limit cycle emerges (Fig. 2C).
To investigate temperature scaling, Arrhenius laws can be assigned to the kinetic rates governing Cyclin B synthesis ($k_{s}$), degradation ($k_{d}$), and Cdk1 activation and inactivation. Even this minimal two-ODE system is sufficient to explain non-Arrhenius period scaling. When individual reactions have different activation energies (Fig. 2D), the ratio $k_{s}/k_{d}$ changes with temperature, shifting the Cyclin B nullcline in the phase plane. If $k_{s}$ increases more steeply with temperature than $k_{d}$, the Cyclin B nullcline moves upward. At high enough temperatures the nullclines cease to intersect on the middle branch of the Cdk1 nullcline, the US disappears, and the limit cycle collapses into a high-Cdk1 fixed point (persistent M phase). Conversely, at low temperatures the system becomes trapped in a low-Cdk1 interphase-like state. These bifurcations generate thermal limits and a curved, non-Arrhenius period–temperature relation (Fig. 2D).
Experimental measurements in cycling Xenopus egg extracts support this mechanistic picture [4]. Within the viable embryonic temperature range, the rising (interphase) and falling (M-phase) segments of the oscillation scale differently with temperature, consistent with Cyclin B synthesis being more temperature-sensitive than its degradation. This difference largely explains the non-Arrhenius scaling of the total period in the physiological range.
@@
Early mechanistic proposals by Hastings and Sweeney suggested that opposing biochemical reactions within the clock could have counterbalancing temperature responses [10, 11]. This idea motivated a series of mathematical models in which different parts of the transcription–translation feedback loop scale differently with temperature [12, 13, 14]. However, perfect compensation of temperature dependencies is not robust to parameter variation or genetic perturbations.
@@
Influence of network structure
The oscillator case studies above already show that temperature responses are shaped by network architecture, not only by the temperature dependence of individual reactions. The deterministic network wiring determines (i) which processes lie on the effective delay/slow manifold and (ii) which parameter combinations enter the period-setting feedback gain. A useful language for organizing these design principles is that of network motifs, which are recurrent small subgraphs such as negative feedback loops and feedforward loops that implement stereotyped dynamical functions in transcriptional and signaling networks [16, 17, 18, 19]. Throughout this subsection we use temperature compensation to mean near-invariance of the oscillation period, and temperature robustness more broadly to include stability of other dynamical features (e.g., amplitude, transient waveforms, or noise sensitivity).
In oscillatory systems, motifs rarely appear in isolation: biologically realistic clocks often combine a delayed negative feedback backbone with additional positive feedback and/or post-translational modification cycles that reshape effective nonlinearity and delay [20, 21]. In particular, a recent systematic comparison of minimal circadian oscillator motifs showed that a pure negative feedback (Goodwin-type) architecture (Fig. 2E) tends to yield the best temperature compensation of the period, whereas introducing positive feedback can improve robustness of the oscillation period to extrinsic noise. Moreover, combining positive and negative feedback can reduce the temperature dependence of noise sensitivity compared to either loop alone [21].
This motif-level tradeoff links naturally to classic control-theoretic perspectives: (i) temperature compensation can emerge from balancing period-increasing and period-decreasing processes (“antagonistic balance”) [22], and (ii) motifs that implement (approximate) integral control can generate adaptation/insensitivity properties that resemble compensation in the appropriate input/output mapping [23]. Complementing these theoretical views, measurements of canonical motifs show that temperature robustness can depend strongly on operating regime: for example, strengthening a negative feedback loop can reduce the temperature sensitivity of transient timing, but at a performance cost (lower expression/dynamic range), and coherent feedforward loop behavior exhibits analogous robustness-performance tradeoffs [24].
In circadian clock models for plants (which are typically multi-loop transcriptional networks), the scanning of broad parameter sets has led to the finding that temperature-robust solutions are enriched for strongly repressive interactions, autoregulation, and three-node feedback loops [25]. This complements motif-by-motif oscillator comparisons by showing that robustness can be associated with specific wiring patterns across large regions of parameter space.
Finally, even when two oscillators share a similar coarse-grained architecture, where time delays and feedback exist in the network can alter the temperature-response outcomes. Both the embryonic cell cycle oscillator (Fig. 2A-D) and circadian models (Fig. 2E-H) have a similar network architecture, combining negative feedback with additional regulatory steps that change the effective feedback strength [26]. However, the consequences for temperature responses differ. In the circadian model, the interconversion between protein states occurs on timescales comparable to the temperature-sensitive synthesis and degradation reactions, allowing redistribution between isoforms to buffer temperature-induced changes and stabilize the period [27]. In the embryonic cell cycle oscillator, by contrast, the dominant slow process is Cyclin B accumulation, whereas the switch between Cdk1 states is very fast. This strong timescale separation means that changes in the relative temperature sensitivities of synthesis and degradation directly shift the nullclines and hence the period, decreasing the internal buffering. Thus, despite similarities in network structure, differences in relative timescales and in how feedback is routed lead to very different temperature responses.
@@
Stochastic network models represent biological processes as proceeding along a complex directed graph where the nodes represent states and the edges transitions. Assigning to each edge a temperature-dependent transition rate, this becomes a continuous time Markov-jump process. For a given process, one can then use the master equation to compute the distribution of first passage times between a given initial and final state [30]. Most importantly, the first order moment of this distribution, the mean first passage time (MFPT), can then be equated to the inverse of the rate at which the process proceeds, while higher-order moments are usually expected to be minimized [31]. This allows to explicitly derive how temperature scaling laws emerge from stochastic network topology.
@@
Recent work has first established this statistical framework for the generic temperature dependence of large biochemical networks which are stochastic and for which each rate follows an Arrhenius law [28]. We note that even at the level of individual reactions, the Arrhenius form derived from Kramers’ theory relies on the assumption of sufficiently large energy barriers. When barriers become comparable to thermal fluctuations, this approximation breaks down and non-Arrhenius scaling can arise already at the single-reaction level, as shown using mean first-passage time approaches [32]. Starting from the master equation of a time-continuous, space-discrete Markov process, the analysis uses the graph-theoretic representation of the MFPT to a target state [33]. The overall rate $r$ is the inverse MFPT, and its temperature dependence $\ln{r(T)}$ can be expanded in a Taylor series whose coefficients are related to the distributions of the total activation energies along the spanning trees $E_{\mathcal{T}}$ and spanning forests of two trees $E_{\mathcal{F}}$ of the stochastic network (see Fig. 3 A) according to:
@@
where $\langle E\rangle$ and $\sigma^{2}$ denote the means and variances of the respective energy distributions. This prediction agrees well with simulations of random networks and published developmental-rate data in flies which have previously been fitted by the phenomenological quadratic exponential described in Part I [34, 35, 36]. This limit requires that the difference in the activation energies along the trees and the forests remains distinct in the large network limit, typically assured by a bias in the activation energies for rates towards the target state. Moreover, it is only fully valid if the network is sufficiently complex to have a macroscopic number of spanning trees and forests.
@@
The master equation approach has also been applied to a specific, and analytically tractable, network topology: a linear chain of $n$ reversible transitions (Fig. 3 C) [29]. Such cascades provide a reasonable representation of many multi-step biochemical processes. However, since they have only one spanning tree they do not fulfill the requirements leading to the result in Eq. (2) [28]. Still, the analysis of this model reveals that, even in this case, the generic quadratic exponential arises. Additionally, it shows how and why this quadratic scaling can break down at extreme temperatures.
Each individual step follows an Arrhenius law with a randomly distributed activation energy. At the reference temperature $T_{0}$, the cascade is assumed to be forward-biased, meaning that forward rates are, on average, larger than backward rates. Computing the MFPT, for example via the technique described in the previous section [33], from the initial to the final state shows that the overall temperature dependence of the rate falls into three regimes.
In the physiologically relevant range around $T_{0}$, the many forward transitions collectively dominate the MFPT. By the law of large numbers, the sum of their activation energies approaches a normal distribution. Just as in Eq. (2) this leads to a quadratic exponential scaling for the overall rate
$\ln r(T)=\langle E\rangle\Delta\beta+\frac{\sigma_{E}^{2}}{2}\Delta\beta^{2}+\ln\left(\frac{n}{\langle 1/r_{f}^{*}\rangle}\right),$ (3)
@@
At low temperatures, however, the system undergoes a sudden transition. As backward reactions become comparably faster, the system spends an increasing amount of time in cycles of forward and backward reactions before reaching the final state. Below a critical temperature $T_{-}$, a single cycle dominates the MFPT. In this regime, the rate scales Arrhenius-like with a large positive activation energy that depends on the sum and difference of the activation energies within this critical cycle.
A similar breakdown of the quadratic exponential occurs at high temperatures. Above a second critical temperature $T_{+}$, one cycle again dominates the MFPT leading to an overall rate that follows the Arrhenius law. However, in this regime the effective activation energy becomes negative. The critical temperatures separating the three regimes are:
@@
Biological processes exhibit diverse temperature responses across scales, yet many of their large-scale features arise from shared physical and organizational principles. In Part I of this review [1], we showed how phenomenological descriptions and microscopic reaction-level theories capture robust regularities in rate–temperature relationships and clarify why deviations from simple Arrhenius behavior are common. Here, in Part II [2], we have focused on how these local temperature dependencies are transformed by network organization into emergent system-level temperature responses.
@@
Declaration of generative AI and AI-assisted technologies in the manuscript preparation process
During the preparation of this work the authors used ChatGPT in order to get feedback on grammar and phrasing. After using this tool, the authors reviewed and edited the content as needed and take full responsibility for the content of the published article.
@@ Acknowledgements
The work is supported by grants from Internal funds KU Leuven (C14/23/130, LG). JBV thanks the German Academic Scholarship Foundation (Studienstiftung des Deutschen Volkes) and USS the Max Planck School Matter to Life, which is funded by the Dieter Schwarz Foundation and the Max Planck Society.